Suppose that you own a 10-acre plot of land that you would like to rent out to wheat farmers. For them, bringing in a harvest involves $35 per acre for seed, $80 per acre for fertilizer, and $70 per acre for equipment rentals and labor. With these inputs, the land will yield 40 bushels of wheat per acre. Now suppose the price at which wheat can be sold is $6 per bushel and that farmers want to earn a normal profit of $10 per acre.

What is the most that any farmer would pay to rent your 10 acres?

- $

What if the price of wheat rose to $7 per bushel?

- $
Answer #1

Total Cost per acre:

TC= seed+fertilizers+equipments and labor= 35+70+80= 185

Revenue: when price= $6

TR=40x6= 240

Profits before rent= 240-185= 55

Economic Profit= 10

Affordable Rent= $45 per acre

If price increases to $7

Total Revenue= 40x7= 280

Profit= 280-185= 95

Economic Profit= 10

Affordable rent= $85 per acre